OpenStudy (anonymous):

∫{dy/y(4y^2-)^(1/2) Find the definite integral

OpenStudy (anonymous):

-1 is under the square root too?

OpenStudy (anonymous):

\[\int\limits_{}^{}{\frac{dy}{y \sqrt{4y^{2}-1}}}\]

OpenStudy (anonymous):

thats the correct way

OpenStudy (anonymous):

I am not sure, thinking

OpenStudy (anonymous):

yeah this is a tricky one...

OpenStudy (anonymous):

I dont know, sorry :(

OpenStudy (anonymous):

I am stuck

OpenStudy (anonymous):

I can help you :D

OpenStudy (anonymous):

Here is the help! :)

OpenStudy (anonymous):

:D Let me write it down and then I'll type it up xP

OpenStudy (anonymous):

ok thank you so much! andras...alright...i'll wait for your response malevolence

OpenStudy (anonymous):

This is the way I did it. I noticed that you have a sqrt in the form of sqrt(u^2-a^2). So rewrite it this way: \[\frac{1}{2}\int\limits \frac{dy}{y \sqrt{y^2-\frac{1}{4}}}\] Use the substitution sec(x)=y so: sec(x)tan(x)dx=dy and tan(x)=sqrt(y-(1/4)) Rewriting then you have: \[\frac{1}{2} \int\limits \frac{\sec(x)\tan(x)dx}{\sec(x)\tan(x)}=\frac{1}{2}\int\limits dx=\frac{x}{2}\] But x=arcsec(y) So the integral is: \[\frac{\sec^{-1}(y)}{2}+C\] However, wolfram did it slightly differently but I'm assuming they are equivalent. http://www.wolframalpha.com/input/?i=integral+of+dx%2F(x(sqrt(4x^2-1)))

OpenStudy (anonymous):

wow..okay...so if we never learn that 1/(u^2+1) is tan inverse? thats the only way to do it? and thanks so much malevolence...

OpenStudy (anonymous):

The trig substitution that I did above is who I know how to do it. The u-sub that wolfram did is crafty. I feel as though they won't be any different however. Try your definite integral and see I suppose.

OpenStudy (anonymous):

And you're welcome :)